If an individual’s preferences are described by the utility function U(X1 , X2 ) = X12 + X22, graph the indifference curve for U = 20 and U= 40.Find the optimal consumption quantities if P1 = US$2.50 ;         P2 = US$ 7.50; and M = US$ 60.


1
Expert's answer
2021-06-17T10:18:34-0400

Utility function

"U(x_1,x_2)=x_1^2+x_2^2"

so,


The indifference curve

"U=20\\V=40"






Given information

"p_1=US\$2.50\\p_2=US\$7.50"


"M=US\$60"


Budget line equation

"p_1x_1+p_2x_2=M..............................................(1)"


substitute these values in the above equation 1

"\$2.50x_1+\$7.50x_2=60....................................(2)\\From\space U(x_1,x_2)=x_1^2+x_2^2\\MRS(x_1,x_2)=\frac{MUx_1}{MUx_2}\\MUx_1=2x_1,MUx_2=2x_2\\MRS(x_1,x_2)=\frac{2x_1}{2x_2}\\=\frac{x_1}{x_2}"


For optimal bundle

equation slope of budget line and MRS

"\frac{p_1}{p_2}=MRS_{(x_1,x_2)}\\"

"\frac{2.5}{7.5}=\frac{x_1}{x_2}"

"x_2(2.5)=(7.5)x_1\\x_2=3x_1....................................................(3)"

putting x2 value in eq 2

we get

"\$2.50x_1+\$7.50x_2=60\\\$2.50x_1+\$7.50\times3x_1=60\\\$2.50x_1+\$22.5x_1=60\\25x_1=60\\x_1=\frac{60}{25}\\x_1=2.4"


putting x1 value in eqn 3

"x_2=3x_1\\=3\times2.4\\x_2=7.2"


x1=2,4

x2=7.2